OpenStudy (anonymous):

how many moles of gass will a 486 ml flask hold at 10 celsius and 500 torr?

OpenStudy (anonymous):

@iPwnBunnies can u help with one more plz

OpenStudy (anonymous):

its gas law

OpenStudy (somy):

use the formula PV = nRT u want to find mole so u meed to make 'n' as subject n = PV/RT P= 500 torr (convert to Pascal) so it'll be around 66 661.2 Pa V= 485 ml (convert to m^3) so it'll be 485/1000 000= 4.85*10^-4 R (is a constant)= 8.31 T= 10 Celsius (convert to Kelvin) so it'll be 10+273= 283 K and now substitute all these into the formula and u'll get the answer

OpenStudy (anonymous):

@arabpride

OpenStudy (anonymous):

@bbcream14

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

oh ok thanks anyway:)

OpenStudy (anonymous):

just do what somy did

OpenStudy (anonymous):

and welcome:)

OpenStudy (anonymous):

@NightMath

OpenStudy (anonymous):

idk if my set up is right?

OpenStudy (anonymous):

Srry I'm not sure about this Question*_*

OpenStudy (anonymous):

thanks for trying:)

OpenStudy (anonymous):

@enano64fs

OpenStudy (anonymous):

Go with @Somy

OpenStudy (anonymous):

is the set up right i got 661.2pa*485000cm^3/8.31g/mol*283k?

OpenStudy (anonymous):

@bashirk

OpenStudy (anonymous):

i just want to knw if my set up is right and which units cancel

OpenStudy (anonymous):

I have no clue how to do this im sorry

OpenStudy (anonymous):

oh ok thanks anyways:)

OpenStudy (anonymous):

@urb0b

OpenStudy (somy):

@Karina1 you have to use SI system units here, that's why you need to convert those units to appropriate units from SI system when you use SI system you don't need to worry about units because you will surely get it right

OpenStudy (somy):

@Karina1 i mean you answer will be with correct unit